OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0

OpenStudy (kirbykirby):

Is that \(\sin^2x+\sin x=0\)? If so, a great big hint is to try substituting \(y=\sin x\), which will give you a "normal" looking quadratic equation. Solve that, then re-substitute back the \(y\)'s with \(\sin x\).

OpenStudy (anonymous):

yes it is that! @kirbykirby

OpenStudy (anonymous):

x = 0, π, 3pi/2 x = 0, π, pi/3, 5pi/3 x = 0, π, pi/3,2pi/3 x = 0, π, 4pi/3, 5pi/3

OpenStudy (anonymous):

those were pi signsand when i did the problem i got C but im not sure if im correct @kirbykirby

OpenStudy (kirbykirby):

What did you get as solutions to the quadratic equation if you substitute \(\sin x\) with another variable?

OpenStudy (anonymous):

to be honest i made a guess but Im not sure how I came up with this conclusion because Ive done the problem once before but I never knew if i got it right lol @kirbykirby

OpenStudy (kirbykirby):

Hm ok well we'll try and go through this together. If you try this substitution: \(y=\sin x\), you end up with \(y^2+y=0\) Factoring gives: \(y(y+1)=0\) So What are the solutions for \(y\)?

OpenStudy (anonymous):

y=0,-1?

OpenStudy (anonymous):

thats what I got when i factored

OpenStudy (kirbykirby):

Yep that's great , now if we substitute back \(y = \sin x \), you get \(\sin x = 0\) and \(\sin x =-1\)

OpenStudy (kirbykirby):

Now you need to find the x-values that satisfy sin x = 0, and sin x = -1

OpenStudy (kirbykirby):

You can use a unit circle for that

OpenStudy (anonymous):

:/ not understanding how to get the solutions @kirbykirby

OpenStudy (kirbykirby):

Well, in the unit circle, you have the coordinates expressed as (cos x, sin x) And the main points of interest on the circle are those that gives values of 0 and -1 for sine: